OpenStudy (anonymous):

solve problem using the fundamental theorem of calculus: f'(x)=e^(-x) and f(5)=1. Find f(2). Please and thank you :)

myininaya (myininaya):

\[f(x)=\int\limits_{}^{}e^{-x} dx+C=-e^{-x}+C\]

myininaya (myininaya):

\[f(5)=-e^{-(5)}+C=1\]

myininaya (myininaya):

so we have enough info to solve for C

myininaya (myininaya):

\[-e^{-5}+C=1 => C=1+e^{-5} => f(x)=-e^{-x}+(1+e^{-5})\]

myininaya (myininaya):

See this is a cal 1 problem not the type of question you posted earlier

myininaya (myininaya):

\[f(2)=-e^{-(2)}+(1+e^{-5})=e^{-5}-e^{-2}+1\]

OpenStudy (anonymous):

oh.....I get it now... It makes complete sense. thank you :)